# A Composable Agentic System for Automated Visual Data Reporting

### v1

@@ Abstract
To address the brittleness of monolithic AI agents, our prototype for automated visual data reporting explores a Human-AI Partnership model. Its hybrid, multi-agent architecture strategically externalizes logic from LLMs to deterministic modules, leveraging the rule-based system Draco [21] for principled visualization design. The system delivers a dual-output: an interactive Observable [17] report with Mosaic [10] for reader exploration, and executable Marimo [1] notebooks for deep, analyst-facing traceability. This granular architecture yields a fully automatic yet auditable and steerable system, charting a path toward a more synergistic partnership between human experts and AI. For reproducibility, our implementation and examples are available at peter-gy.github.io/VISxGenAI-2025.
@@ Introduction
This partnership model is motivated by a core challenge in data analysis. Even a technically flawless agent may fail if it perfectly executes the wrong task. Human intent is often ambiguous at the outset and evolves as insights emerge. A minor misalignment can render an agent’s entire output useless, forcing costly and unstable regeneration that erodes user trust and agency. Genuine human–AI partnership requires systems explicitly designed to address this challenge: providing auditable, high-quality outputs while enabling domain experts to iteratively refine them. To investigate this model, we developed a prototype agentic system for automated visual data reporting [^1] Our workshop challenge submission is available online.. The design of our implementation was guided by four high-level principles, adapting established guidelines for human-AI interaction [2] to prioritize human agency alongside automation [9].
@@ 1 System Design and Implementation
Our prototype is a multi-agent system implemented using DSPy [12]. As A Composable Agentic System for Automated Visual Data Reporting shows, the main Orchestrator  A coordinates a workflow of specialized, hybrid agents that combine LLM reasoning with deterministic logic. It accepts as input a dataset URI and high-level user intent, such as the overall goal and desired length of the report. To ensure full audibility, the Orchestrator collects and stores detailed execution traces of all agent interactions in Langfuse [13], an external observability service.
@@
2. Analysis and Materialization. With a clean dataset, the Insight Planner  F operates within a Reasoning and Acting (ReAct) loop [19], using a dedicated tool to retrieve hints from the statistical profile and ground its planning in data. For each planned insight, the Dataset Deriver  G then drafts and, if necessary, repairs a DuckDB SQL query [18], also outputting semantic roles for the generated fields. Adhering to our Granularity principle, the programmatic Dataset Publisher  H executes each query and stores the resulting materialized dataset as a versioned Apache Parquet file in Cloudflare R2, an S3-compatible object store [3, 5].
@@
3. Visualization. We avoid monolithic LLM-based chart design, a token-intensive approach whose recommendations only moderately align with human preferences and remain widely inconsistent [20]. Instead, we use LLM-generated metadata from prior stages to inform a deterministic visualization process. The semantic roles from the Dataset Deriver  G and the analytical task from the Insight Planner  F are used to programmatically construct a partial Draco specification that defines what to visualize [21]. The Dataset Visualizer  I then delegates the task of determining how to visualize it to Draco’s rule-based solver. Based on its computational analysis of the data’s properties (e.g., cardinality, skew), Draco can recommend complex, appropriate charts like a faceted scatterplot with a symmetric logarithmic scale–a nuanced design decision difficult for a pure LLM approach. This provides another advantage: if a recommendation is suboptimal, its logic is transparent. We can systematically improve future results by adjusting Draco’s knowledge base, a key benefit of a white-box system.
@@
4. Reporting and Publication. The final stage packages all artifacts for human interaction. A vision-language model, coordinated by Report Narrator  J , generates textual descriptions for each base64-encoded chart image and structures them into a cohesive narrative. The Dataset Reporter  K then assembles these with granular artifacts produced by upstream agents. Because all these components are decoupled from presentation, they are modality-agnostic and can be composed into diverse formats, from an interactive web report to a static PDF. Our system demonstrates this flexibility by programmatically constructing a dual output from these same artifacts: JavaScript source for an interactive Observable 2.0 report [16], which a self-hosted Node.js service [15, 17, 8] builds into a static website, and a Python-based executable Marimo notebook [1] for deep, analyst-facing traceability. The Observable report uses Mosaic [10] to link each visualization with its Quak data table [14], an approach inspired by narrative visualization [11] that combines the AI’s authored insights with interactive reader-driven exploration. The Marimo notebook, in turn, serves as an executable trace for a single insight, revealing the precise data transformation and visualization logic. This dual-output system provides distinct entry points for expert interaction and targeted modification.
@@
By deconstructing the workflow into well-scoped tasks, our modular design allows for a “right-size” model strategy [4]. Smaller, faster models like Gemini 2.5 Flash [6] handle the majority of tasks, reducing overall latency, while the more powerful Gemini 2.5 Pro [7] is reserved for the most demanding functions: generating complex SQL  G and authoring descriptive visual narratives  J . This design makes such systems more future-ready, allowing models to be swapped individually and supporting the use of open-weight, self-hosted alternatives.
@@ 2 Discussion
Testing on eight diverse datasets demonstrates our system’s generalizability, producing meaningful reports without any domain-specific tuning [^2] All discussed reports are available for exploration here. . The value of our Human-AI Partnership model is best illustrated by the distinct workflows it enables: systemic steerability for the analyst creating the report, and open-ended exploration for the reader consuming it. Steerability for the Analyst. An early report featured a weak insight, faceting a scatterplot by the highly skewed GraphicsReplicabilityStamp field2. Instead of fragile prompt engineering, our architecture supports a durable fix. Using execution traces, an analyst can pinpoint that the Insight Planner  F was misled by a hint from its Dataset Profile Query Tool. The analyst then intervenes by modifying the tool’s logic, adding a rule to only suggest fields for segmentation that exceed a statistical diversity threshold. This dataset-agnostic refinement permanently improves the system’s knowledge base, preventing similar flawed insights in all future runs. Explorability for the Reader. The partnership extends to the reader. Our reports use Mosaic [10] to enable exploration beyond the AI’s authored narrative. A static chart of downloads vs. citations is enhanced with a linked data table; readers can answer complex ad-hoc questions (e.g., “What about VIS papers in the last five years?”) by cross-filtering with attributes like conference and year from the linked data table, which are not encoded in the main visualization.

### v2

@@ Abstract
To address the brittleness of monolithic AI agents, our prototype for automated visual data reporting explores a Human-AI Partnership model. Its hybrid, multi-agent architecture strategically externalizes logic from LLMs to deterministic modules, leveraging the rule-based system Draco [Yang:2023:Draco2] for principled visualization design. The system delivers a dual-output: an interactive Observable [Observable:2025:NBKit] report with Mosaic [Heer:2024:Mosaic] for reader exploration, and executable Marimo [Marimo:2023:Notebooks] notebooks for deep, analyst-facing traceability. This granular architecture yields a fully automatic yet auditable and steerable system, charting a path toward a more synergistic partnership between human experts and AI. For reproducibility, our implementation and examples are available at peter-gy.github.io/VISxGenAI-2025.
@@ Introduction
This partnership model is motivated by a core challenge in data analysis. Even a technically flawless agent may fail if it perfectly executes the wrong task. Human intent is often ambiguous at the outset and evolves as insights emerge. A minor misalignment can render an agent’s entire output useless, forcing costly and unstable regeneration that erodes user trust and agency. Genuine human–AI partnership requires systems explicitly designed to address this challenge: providing auditable, high-quality outputs while enabling domain experts to iteratively refine them. To investigate this model, we developed a prototype agentic system for automated visual data reporting [^1] Our workshop challenge submission is available online.. The design of our implementation was guided by four high-level principles, adapting established guidelines for human-AI interaction [Amershi:2019:HAIGuidelines] to prioritize human agency alongside automation [Heer:2019:AgencyPlus].
@@ 1 System Design and Implementation
Our prototype is a multi-agent system implemented using DSPy [Khattab:2023:DSPy]. As A Composable Agentic System for Automated Visual Data Reporting shows, the main Orchestrator  A coordinates a workflow of specialized, hybrid agents that combine LLM reasoning with deterministic logic. It accepts as input a dataset URI and high-level user intent, such as the overall goal and desired length of the report. To ensure full audibility, the Orchestrator collects and stores detailed execution traces of all agent interactions in Langfuse [Langfuse:2025:LLMPlatform], an external observability service.
@@
2. Analysis and Materialization. With a clean dataset, the Insight Planner  F operates within a Reasoning and Acting (ReAct) loop [Yao:2023:ReAct], using a dedicated tool to retrieve hints from the statistical profile and ground its planning in data. For each planned insight, the Dataset Deriver  G then drafts and, if necessary, repairs a DuckDB SQL query [Raasveldt:2019:DuckDB], also outputting semantic roles for the generated fields. Adhering to our Granularity principle, the programmatic Dataset Publisher  H executes each query and stores the resulting materialized dataset as a versioned Apache Parquet file in Cloudflare R2, an S3-compatible object store [ApacheParquet, CloudflareR2].
@@
3. Visualization. We avoid monolithic LLM-based chart design, a token-intensive approach whose recommendations only moderately align with human preferences and remain widely inconsistent [Wang:2025:DracoGPT]. Instead, we use LLM-generated metadata from prior stages to inform a deterministic visualization process. The semantic roles from the Dataset Deriver  G and the analytical task from the Insight Planner  F are used to programmatically construct a partial Draco specification that defines what to visualize [Yang:2023:Draco2]. The Dataset Visualizer  I then delegates the task of determining how to visualize it to Draco’s rule-based solver. Based on its computational analysis of the data’s properties (e.g., cardinality, skew), Draco can recommend complex, appropriate charts like a faceted scatterplot with a symmetric logarithmic scale–a nuanced design decision difficult for a pure LLM approach. This provides another advantage: if a recommendation is suboptimal, its logic is transparent. We can systematically improve future results by adjusting Draco’s knowledge base, a key benefit of a white-box system.
@@
4. Reporting and Publication. The final stage packages all artifacts for human interaction. A vision-language model, coordinated by Report Narrator  J , generates textual descriptions for each base64-encoded chart image and structures them into a cohesive narrative. The Dataset Reporter  K then assembles these with granular artifacts produced by upstream agents. Because all these components are decoupled from presentation, they are modality-agnostic and can be composed into diverse formats, from an interactive web report to a static PDF. Our system demonstrates this flexibility by programmatically constructing a dual output from these same artifacts: JavaScript source for an interactive Observable 2.0 report [Observable:2025:NB2], which a self-hosted Node.js service [NodeJS, Observable:2025:NBKit, Gyarmati:2025:NBBuilder] builds into a static website, and a Python-based executable Marimo notebook [Marimo:2023:Notebooks] for deep, analyst-facing traceability. The Observable report uses Mosaic [Heer:2024:Mosaic] to link each visualization with its Quak data table [Manz:2025:Quak], an approach inspired by narrative visualization [Heer:2010:NarrativeVis] that combines the AI’s authored insights with interactive reader-driven exploration. The Marimo notebook, in turn, serves as an executable trace for a single insight, revealing the precise data transformation and visualization logic. This dual-output system provides distinct entry points for expert interaction and targeted modification.
@@
By deconstructing the workflow into well-scoped tasks, our modular design allows for a “right-size” model strategy [Belcak:2025:SLM]. Smaller, faster models like Gemini 2.5 Flash [Gemini25Flash] handle the majority of tasks, reducing overall latency, while the more powerful Gemini 2.5 Pro [Gemini25Pro] is reserved for the most demanding functions: generating complex SQL  G and authoring descriptive visual narratives  J . This design makes such systems more future-ready, allowing models to be swapped individually and supporting the use of open-weight, self-hosted alternatives.
@@ 2 Discussion
Testing on eight diverse datasets demonstrates our system’s generalizability, producing meaningful reports without any domain-specific tuning [^2] All discussed reports are available for exploration here. . The value of our Human-AI Partnership model is best illustrated by the distinct workflows it enables: systemic steerability for the analyst creating the report, and open-ended exploration for the reader consuming it. Steerability for the Analyst. An early report featured a weak insight, faceting a scatterplot by the highly skewed GraphicsReplicabilityStamp field2. Instead of fragile prompt engineering, our architecture supports a durable fix. Using execution traces, an analyst can pinpoint that the Insight Planner  F was misled by a hint from its Dataset Profile Query Tool. The analyst then intervenes by modifying the tool’s logic, adding a rule to only suggest fields for segmentation that exceed a statistical diversity threshold. This dataset-agnostic refinement permanently improves the system’s knowledge base, preventing similar flawed insights in all future runs. Explorability for the Reader. The partnership extends to the reader. Our reports use Mosaic [Heer:2024:Mosaic] to enable exploration beyond the AI’s authored narrative. A static chart of downloads vs. citations is enhanced with a linked data table; readers can answer complex ad-hoc questions (e.g., “What about VIS papers in the last five years?”) by cross-filtering with attributes like conference and year from the linked data table, which are not encoded in the main visualization.